OpenStudy (anonymous):

Find the 18th term in the sequence. 9, 5, 1, -3, ...

Mehek (mehek14):

each term is decreasing by 4

Mehek (mehek14):

9 - 4 = 5 5 - 4 = 1 1 - 4 = -3

OpenStudy (anonymous):

(-4(x-1))+9 where x is the term in the sequence. So..-59